The most common isotope of copper is copper-63 with a measured mass of 62.929601u.

  1. Determine the mass defect;
  2. Determine the binding energy;
  3. Determine the binding energy per nucleon.
1
Expert's answer
2021-06-03T09:16:52-0400

1.

m(proton) = 1.007276 u

m(neutron) = 1.008665 u

Cu-63 has 29 protons and 34 neutrons.

m(defect) = -(62.929601 - 1.007276"\times"29 - 1.008665"\times"34) = 0.576013 u

2.

0.576013 u = 1.66054"\times"10-27 kg

E(per atom) = mc2 = 1.66054"\times"10-27 "\times" 299,792,4582 = 1.492419"\times"10-10 J

3.

E(per nuclon) = "\frac{1.492419\times10^{-10} J }{63} = 2.368919\times 10^{-12} J"